# Development & Characterization of Electrodes for large-scale Xenon Time Projection Chambers

### v1

@@ 4.2.2 Argon Calibration
Breakdown voltage or HV instability depends, among other factors, on the gas composition of the environment as well as the water content in it [humidity, humidity_2, high_voltage_book:1991]. Therefore, the humidity inside the test box was monitored by a humidity sensor. It was also used as a proxy for the air concentration as the GAr was flushed into the box. Note that the breakdown voltage in argon gas is lower than that in air at our operating conditions [Raizer_book:1991]; therefore, the humidity should correlate with the breakdown voltage in our setup. To that end, a calibration run was performed to verify the correlation between GAr flush time, humidity, and the HV instability of the setup.
@@
In this calibration run, most ramps have a ramp-up speed of 500 V/s. At the beginning of the calibration campaign (negative times in Figure 14), many ramps were performed to condition the system, removing the remaining dust and impurities [lucie:2019, conditioning:1982], resulting in a gradual increase in $V_{c}$. At time =0=0, the GAr valve was opened and both the humidity and $V_{c}$ were reduced. Since there was an overnight break during the test, the argon escaped from the box, causing $V_{c}$ to increase. Therefore, an extra flush time is required to bring $V_{c}$ back to the level it was before the pause. The dashed line in Figure 14 marks the time window that was eliminated. The argon flow was closed occasionally to confirm that the change in GAr concentration caused the reduction in $V_{c}$. Those periods were also removed from the plot such that we obtained the effective flush time.
The $V_{c}$ and humidity profiles were correlated throughout the test, the latter being limited by the sensitivity of the sensor. From the result, a flush time at 20 NL/min of slightly more than $2\text{\,}\mathrm{h}$ resulted in a relatively stable $V_{c}$. This implied that the uncertainty of the GAr concentration on the HV instability became subdominant. From $1.5\text{\,}\mathrm{h}$ on until the end of the measurement, the relative change in $V_{c}$ was 16.4%. The absolute breakdown voltage $|V_{c}|$ plateaued at $18.0\text{\,}\mathrm{k}\mathrm{V}$ after $2.26\text{\,}\mathrm{h}$ of flushing, corresponding to a bulk field of $5.1\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$. The breakdown voltage between polished spheres in gaseous argon parameterized by the PyBoltz-Townsend model, at a pressure of $1\text{\,}\mathrm{a}\mathrm{t}\mathrm{m}$ and a gap distance of $3.5\text{\,}\mathrm{c}\mathrm{m}$, was reported at $21.7\text{\,}\mathrm{k}\mathrm{V}$ [Ar_Xe_breakown:Norman2022], resulting in a bulk field of $6.2\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$. The obtained lower value from our measurement is expected, as the surface of the plates was not mirror-polished.
@@ 4.2.3 Imaging High-Voltage Breakdown
As shown in the upper panel of Figure 15, the HV was increased in steps until a breakdown occurred. The mean brightness of each frame captured by the third camera, was used to correlate the light emission to the current profile in Figure 15 (lower left). We computed the average brightness value of all pixels per frame (gray line) and the average value of all pixels in a $50\times 25$ pixel region around the broken leg (blue line). A gradual increase in the mean brightness was observed at the location of the broken leg, which was directly correlated with the increase in the current profile prior to the breakdown. At breakdown the faint glow developed into a continuous arc between the location of the broken hex leg and the ground plate. The same increase in glow was also observed from the other two cameras that were mounted above the setup with long-exposure images. This basic test indicated that it was possible to utilize the setup to identify faint glow associated with field emission effects and correlate them with the current profile and specific features on the electrode mesh.
@@ 4.3 High-Voltage Performance
Data that were collected throughout the HV tests listed in Table 3 were used to assess the performance of the mesh electrode. Tests that were performed after stage 1 repair, resulted in 19 breakdowns, out of which 16 occurred along the stretching ring or on the electrode frame, and only three on the mesh itself which are shown in Figure 16. During HV testing that followed stage 2, only three breakdowns were observed, all on the electrode frame. The breakdowns at the stretching ring could be potentially attributed to the presence of sharp corners in its geometry as well as lack of electropolishing. However, as described in subsection 3.2, this ring would be removed after the final stretching and installation of the mesh on the electrode frame. Breakdowns on the frame could result from sharp edges of the screw heads which were not yet covered as in the final installation. In addition, the aluminum ground plate with its edges not perfectly rounded ended just below the frame. Breakdowns that occurred on the mesh in stage 1 could have resulted from the presence of dust particles as the tests were performed in a non-cleanroom environment. The breakdowns that were observed on the electrode mesh were neither correlated with the locations of known features nor with the locations of features that were repaired via mechanical treatment of laser welding. Test results described above indicate that both the welding seam and the laser welded repair locations could withstand HV without causing detrimental effects such as field emission or breakdown.
@@
To quantify the performance of the electrode the obtained breakdown data were fitted with a three parameter Weibull distribution with the use of the Reliability software framework [matthew_reid_2020_3938000]. The Weibull distribution is suitable for analyzing phenomena whose probability of occurrence below a certain threshold is close to zero [HV_engineering_Faruk]. Hence, it is commonly employed in reliability analysis in electrical and mechanical engineering to evaluate components and system performance [reliability_modarres]. The fit results as well as the empirical cumulative distribution function (ECDF) obtained from the data are shown in Figure 17. From the obtained fit results one can ascertain a bulk field of $3.1\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ at 95% survival probability in GAr, with upper and lower bounds of $2.8\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ and $3.4\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$, respectively.
@@
Figure 17: Survival probability that was obtained using all suitable data acquired throughout the HV testing of the cathode electrode. Purple squares depict the values of empirical cumulative distribution function (ECDF) of the calculated bulk fields at breakdown, while the survival function that was obtained from the fitted three parameter Weibull distribution is shown in black. The ECDF is plotted as 1-ECDF, and right-censored data are not shown. The dashed lines correspond to the lower (left) and upper (right) bounds on the $3.1\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ field at 95% survival probability (middle). The gray band corresponds to the 95% confidence bounds on the bulk field at breakdown obtained with the use of the Reliability [matthew_reid_2020_3938000] fitting framework.
@@
Subsequently, we can infer the performance of the electrode in GXe and LXe. We take the conditions in our setup (GAr, 1.0 bar, $4.97\text{\,}\mathrm{c}\mathrm{m}$ gap distance, and $293\text{\,}\mathrm{K}$) and typical values in a dual-phase xenon TPC (GXe, 1.8 bar, $5.4\text{\,}\mathrm{c}\mathrm{m}$ gap distance, and $177\text{\,}\mathrm{K}$). Applying Paschen’s curves [Ar_Xe_breakown_temp:Massarczyk2016, Ar_Xe_breakown:Norman2022] for the case of uniform fields in both conditions results in a ratio of 1.6 and 2.4 from GAr to GXe.
@@
In a second step, we extrapolate the values from GXe to LXe. Although the ionization threshold of LXe was found at around $700\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ [LXe_thresholds_Aprile:2014], breakdowns were recorded at fields lower than this threshold. In a configuration of large electrodes a breakdown field of less than $100\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ was reported [lz_electrodes, xebra:2019]. Various theories have been proposed to account for this difference, such as an earlier breakdown due to the presence of GXe bubbles [LXe_ionization_bubble:1994]. A field enhancement of a factor 1.2 inside such a bubble cavity is calculated, using the dielectric sphere equation [high_voltage_book:1991] with the dielectric constant of 1.88 for LXe [dielectric_constant_CRC:2017, dielectric_constant:1964]. Combining this enhancement factor with the conversion factors from GAr to GXe, the conversion from GAr to LXe is at least a factor of 1.9. As described, the breakdown field at 95% survival rate is at $3.1\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ from our measurement. With the minimum conversion factor of 1.9, the hexagonal electrode should withstand a field strength of $5.9\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ in LXe.
@@ 5 Conclusions & Outlook
In this work we have developed and showcased the capabilities to produce parallel wire and mesh electrodes from simulations and mechanical design to subsequent manufacturing, assembly and HV testing. For wire electrodes, we have developed a new and improved wire installation technique and successfully utilized it to wire a fully functional electrode. Repairing features in a hexagonal mesh electrode by laser-welding yielded a structurally stable mesh that did not exhibit any observable HV issues. Furthermore, a HV test setup was constructed in which we have then tested a hexagonal mesh electrode in GAr atmosphere, reaching field strengths up to $\sim$$4\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$. Having passed successfully the HV tests in GAr as well as further tests in LXe at the PANCAKE facility [pancake], the parallel wire grid and the mesh electrode described in this work were installed in the XENONnT TPC as a replacement of the anode and the cathode, respectively.
The next generation of multi-ton xenon-based TPCs such as XLZD [XLZD:2024gxx] will require new approaches when it comes to design, production and assembly of electrodes and other detector components that span several meters in diameter. The work presented here is part of a comprehensive R&D program aimed at constructing and testing detector components for the current and next generation of TPCs conducted at Karlsruhe Institute of Technology and the University of Freiburg. At the University of Freiburg the aforementioned PANCAKE test platform, measuring $2.75\text{\,}\mathrm{m}$ in diameter, was successfully developed and commissioned as a test facility for full scale detector components. Concurrently, at the Karlsruhe Institute of Technology we pursue a multifaceted R&D program aimed at developing large-scale electrodes. Small, centimeter scale samples of electrodes are being studied at the bite-sized High Voltage setup for Electrodes (bHiVE) in GAr, air and vacuum [Simon2024]. These studies are aimed at exploring ways for enhancing the HV performance of electrodes via coating, chemical treatment and cleaning techniques, which could be then applied to meter scale electrodes.
Furthermore, fully-automated solutions for both installation and assembly of meter scale electrode grids and their subsequent HV testing are being explored. A one-meter scale prototype of a robotic system with micrometer precision for HV assessment of electrodes was successfully developed and commissioned. This High-voltage Coordinate Unit for Targeted Inspection of Electrodes (HiCUTIE) is currently being upgraded [Anna2025] to facilitate a fully automated HV assessment of large-scale electrode performance in cleanroom conditions.
@@ Acknowledgments
We thank the technical staff members from local workshops at KIT and University of Freiburg for providing excellent technical solutions in a very cooperative environment. This work is supported in part by the German Federal Ministry of Research, Technology and Space through grants 05A23VK3, 05A23VF1 and 05A23UM1 in the ErUM-Pro funding line, by the German Research Foundation (DFG) through INST 39/1095-1 FUGG, the Structure and Innovation Fund of the state of Baden-Württemberg (SI-BW), and by the Helmholtz Initiative and Networking Fund through grant W2/W3-118. A.D. was supported by the Cluster of Excellence “Precision Physics, Fundamental Interactions, and Structure of Matter” (PRISMA+ EXC 2118/1) funded by DFG within the German Excellence Strategy (Project ID 390831469). We appreciate support through the graduate school KSETA at KIT. We also gratefully acknowledge support from Kaixuan Ni (University of California San Diego) and Laura Baudis (University of Zürich) as well as helpful discussions and exchanges with Jacques Pienaar (Weizmann Institute of Science) and Hans-Werner Ortjohann (University of Münster).

### v2

@@ 4.2.2 Argon Calibration
Breakdown voltage or HV instability depends, among other factors, on the gas composition of the environment as well as the water content in it [55, 71, 53]. Therefore, the humidity inside the test box was monitored by a humidity sensor. It was also used as a proxy for the air concentration as the GAr was flushed into the box. Note that the breakdown voltage in argon gas is lower than that in air at our operating conditions [72]; therefore, the humidity should correlate with the breakdown voltage in our setup. To that end, a calibration run was performed to verify the correlation between GAr flush time, humidity, and the HV instability of the setup.
@@
In this calibration run, most ramps have a ramp-up speed of 500 V/s. At the beginning of the calibration campaign (negative times in Figure 14), many ramps were performed to condition the system, removing the remaining dust and impurities [83, 67], resulting in a gradual increase in $V_{c}$. At time =0=0, the GAr valve was opened and both the humidity and $V_{c}$ were reduced. Since there was an overnight break during the test, the argon escaped from the box, causing $V_{c}$ to increase. Therefore, an extra flush time is required to bring $V_{c}$ back to the level it was before the pause. The dashed line in Figure 14 marks the time window that was eliminated. The argon flow was closed occasionally to confirm that the change in GAr concentration caused the reduction in $V_{c}$. Those periods were also removed from the plot such that we obtained the effective flush time.
The $V_{c}$ and humidity profiles were correlated throughout the test, the latter being limited by the sensitivity of the sensor. From the result, a flush time at 20 NL/min of slightly more than $2\text{\,}\mathrm{h}$ resulted in a relatively stable $V_{c}$. This implied that the uncertainty of the GAr concentration on the HV instability became subdominant. From $1.5\text{\,}\mathrm{h}$ on until the end of the measurement, the relative change in $V_{c}$ was 16.4%. The absolute breakdown voltage $|V_{c}|$ plateaued at $18.0\text{\,}\mathrm{k}\mathrm{V}$ after $2.26\text{\,}\mathrm{h}$ of flushing, corresponding to a bulk field of $5.1\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$. The breakdown voltage between polished spheres in gaseous argon parameterized by the PyBoltz-Townsend model, at a pressure of $1\text{\,}\mathrm{a}\mathrm{t}\mathrm{m}$ and a gap distance of $3.5\text{\,}\mathrm{c}\mathrm{m}$, was reported at $21.7\text{\,}\mathrm{k}\mathrm{V}$ [66], resulting in a bulk field of $6.2\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$. The obtained lower value from our measurement is expected, as the surface of the plates was not mirror-polished.
@@ 4.2.3 Imaging High-Voltage Breakdown
As shown in the upper panel of Figure 15, the HV was increased in steps until a breakdown occurred. The mean brightness of each frame captured by the third camera was used to correlate the light emission to the current profile in Figure 15 (lower left). We computed the average brightness value of all pixels per frame (gray line) and the average value of all pixels in a $50\times 25$ pixel region around the broken leg (blue line). A gradual increase in the mean brightness was observed at the location of the broken leg, which was directly correlated with the increase in the current profile prior to the breakdown. At breakdown the faint glow developed into a continuous arc between the location of the broken hex leg and the ground plate. The same increase in glow was also observed from the other two cameras that were mounted above the setup with long-exposure images. This basic test indicated that it was possible to utilize the setup to identify faint glow associated with field emission effects and correlate them with the current profile and specific features on the electrode mesh.
@@ 4.3 High-Voltage Performance
Data that were collected throughout the HV tests listed in Table 5 were used to assess the performance of the mesh electrode. Tests that were performed after stage 1 repair resulted in 19 breakdowns, out of which 16 occurred along the stretching ring or on the electrode frame, and only three on the mesh itself which are shown in Figure 16. During HV testing that followed stage 2, only three breakdowns were observed, all on the electrode frame. The breakdowns at the stretching ring could be potentially attributed to the presence of sharp corners in its geometry as well as lack of electropolishing. However, as described in subsection 3.2, this ring would be removed after the final stretching and installation of the mesh on the electrode frame. Breakdowns on the frame could result from sharp edges of the screw heads which were not yet covered as in the final installation. In addition, the aluminum ground plate with its edges not perfectly rounded ended just below the frame. Breakdowns that occurred on the mesh in stage 1 could have resulted from the presence of dust particles as the tests were performed in a non-cleanroom environment. The breakdowns that were observed on the electrode mesh were neither correlated with the locations of known features nor with the locations of features that were repaired via mechanical treatment of laser welding. Test results described above indicate that both the welding seam and the laser welded repair locations could withstand HV without causing detrimental effects such as field emission or breakdown.
@@
To quantify the performance of the electrode the obtained breakdown data were fitted with a three parameter Weibull distribution with the use of the Reliability software framework [73]. The Weibull distribution is suitable for analyzing phenomena whose probability of occurrence below a certain threshold is close to zero [75]. Hence, it is commonly employed in reliability analysis in electrical and mechanical engineering to evaluate components and system performance [62]. The fit results as well as the empirical cumulative distribution function (ECDF) obtained from the data are shown in Figure 17. From the obtained fit results one can ascertain a bulk field of $3.1\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ at 95% survival probability in GAr, with upper and lower bounds of $2.8\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ and $3.4\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$, respectively.
@@
Figure 17: Survival probability that was obtained using all suitable data acquired throughout the HV testing of the cathode electrode. Purple squares depict the values of empirical cumulative distribution function (ECDF) of the calculated bulk fields at breakdown, while the survival function that was obtained from the fitted three parameter Weibull distribution is shown in black. The ECDF is plotted as 1-ECDF, and right-censored data are not shown. The dashed lines correspond to the lower (left) and upper (right) bounds on the $3.1\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ field at 95% survival probability (middle). The gray band corresponds to the 95% confidence bounds on the bulk field at breakdown obtained with the use of the Reliability [73] fitting framework.
@@
Subsequently, we can infer the performance of the electrode in GXe and LXe. We take the conditions in our setup (GAr, 1.0 bar, $4.97\text{\,}\mathrm{c}\mathrm{m}$ gap distance, and $293\text{\,}\mathrm{K}$) and typical values in a dual-phase xenon TPC (GXe, 1.8 bar, $5.4\text{\,}\mathrm{c}\mathrm{m}$ gap distance, and $177\text{\,}\mathrm{K}$). Applying Paschen’s curves [59, 66] for the case of uniform fields in both conditions results in a ratio of 1.6 and 2.4 from GAr to GXe.
@@
In a second step, we extrapolate the values from GXe to LXe. Although the ionization threshold of LXe was found at around $700\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ [18], breakdowns were recorded at fields lower than this threshold. In a configuration of large electrodes a breakdown field of less than $100\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ was reported [56, 82]. Various theories have been proposed to account for this difference, such as an earlier breakdown due to the presence of GXe bubbles [51]. A field enhancement of a factor 1.2 inside such a bubble cavity is calculated, using the dielectric sphere equation [53] with the dielectric constant of 1.88 for LXe [37, 14]. Combining this enhancement factor with the conversion factors from GAr to GXe, the conversion from GAr to LXe is at least a factor of 1.9. As described, the breakdown field at 95% survival rate is at $3.1\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ from our measurement. With the minimum conversion factor of 1.9, the hexagonal electrode should withstand a field strength of $5.9\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$ in LXe.
@@ 5 Conclusions & Outlook
In this work we have developed and showcased the capabilities to produce parallel wire and mesh electrodes from simulations and mechanical design to subsequent manufacturing, assembly and HV testing. For wire electrodes, we have developed a new and improved wire installation technique and successfully utilized it to wire a fully functional electrode. Repairing features in a hexagonal mesh electrode by laser-welding yielded a structurally stable mesh that did not exhibit any observable HV issues. Furthermore, a HV test setup was constructed in which we have then tested a hexagonal mesh electrode in GAr atmosphere, reaching field strengths up to $\sim$$4\text{\,}\mathrm{k}\mathrm{V}\mathrm{/}\mathrm{c}\mathrm{m}$. Having passed successfully the HV tests in GAr as well as further tests in LXe at the PANCAKE facility [31], the parallel wire grid and the mesh electrode described in this work were installed in the XENONnT TPC as a replacement of the anode and the cathode, respectively.
The next generation of multi-ton xenon-based TPCs such as XLZD [3] will require new approaches when it comes to design, production and assembly of electrodes and other detector components that span several meters in diameter. The work presented here is part of a comprehensive R&D program aimed at constructing and testing detector components for the current and next generation of TPCs conducted at Karlsruhe Institute of Technology and the University of Freiburg. At the University of Freiburg the aforementioned PANCAKE test platform, measuring $2.75\text{\,}\mathrm{m}$ in diameter, was successfully developed and commissioned as a test facility for full scale detector components. Concurrently, at the Karlsruhe Institute of Technology we pursue a multifaceted R&D program aimed at developing large-scale electrodes. Small, centimeter scale samples of electrodes are being studied at the bite-sized High Voltage setup for Electrodes (bHiVE) in GAr, air and vacuum [44]. These studies are aimed at exploring ways for enhancing the HV performance of electrodes via coating, chemical treatment and cleaning techniques, which could be then applied to meter scale electrodes.
Furthermore, fully-automated solutions for both installation and assembly of meter scale electrode grids and their subsequent HV testing are being explored. A one-meter scale prototype of a robotic system with micrometer precision for HV assessment of electrodes was successfully developed and commissioned. This High-voltage Coordinate Unit for Targeted Inspection of Electrodes (HiCUTIE) is currently being upgraded [46] to facilitate a fully automated HV assessment of large-scale electrode performance in cleanroom conditions.
@@ Acknowledgments
We thank the technical staff members from local workshops at KIT and University of Freiburg for providing excellent technical solutions in a very cooperative environment. This work is supported in part by the German Federal Ministry of Research, Technology and Space through grants 05A23VK3, 05A23VF1 and 05A23UM1 in the ErUM-Pro funding line, by the German Research Foundation (DFG) through INST 39/1095-1 FUGG, the Structure and Innovation Fund of the state of Baden-Württemberg (SI-BW), by the Helmholtz Initiative and Networking Fund through grant W2/W3-118 and by the U.S. National Science Foundation grant 2112802. A.D. was supported by the Cluster of Excellence “Precision Physics, Fundamental Interactions, and Structure of Matter” (PRISMA+ EXC 2118/1) funded by DFG within the German Excellence Strategy (Project ID 390831469). We appreciate support through the graduate school KSETA at KIT. We also gratefully acknowledge support from Laura Baudis (University of Zürich), as well as helpful discussions and exchanges with Jacques Pienaar (Weizmann Institute of Science) and Hans-Werner Ortjohann (University of Münster).